The guitar string wire in item 2 was 1.60 m long with a diameter of 0.20 cm. Approximately what tension force would break it?


1
Expert's answer
2021-11-08T09:11:51-0500

The tension depends on the strain and young's modulus for the string:


"F=A[\sigma]=A[\epsilon E]=A\frac{\Delta L}{L}E=\frac{\pi d^2}4·\frac{\Delta L}{L}·E,\\\space\\
F=1.96·10^{-6}\Delta LE\text{ N}."